OpenStudy (anonymous):

A car averages 40 miles per hour for the first 6 hours of a trip and averages 60 miles per hour for each additional hour of travel time. If the average speed for the entire trip is 55 miles per hour, how many hours long is the trip?

OpenStudy (anonymous):

come on some one please help

OpenStudy (anonymous):

ok absolutely the time is beyond 6 hours so let the additional hours is x you can get (40*6+60*x)/(6+x)=55 ->x=18

OpenStudy (anonymous):

But the answer given is 24 hrs ... I don't know the methodology

OpenStudy (anonymous):

my answer just the beyond hours but the totle time is 6+x=6+18=24

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Wait, I'm confused. Why is it 24?